Find the second Taylor polynomial of

f(x,y) = xy+3y^2-2 at (1,2)


1
Expert's answer
2021-05-11T08:58:56-0400

Second Taylor degree polynomial can be written as -


"Q(x,y)" = "L(x,y)+\dfrac{fxx_{(1},_{2)}}{2}{(x-1)}^{2}+ {fxy_{(1},_{2)}}{(x-1)}{(y-2)}+\dfrac{fyy_{(1},_{2)}}{2}(y-2)^{2}.......A)"


"L(X,Y)=f(1,2)\ +fx(1,2)(x-1)+fy(1,2)(y-2).........B)"


"f(x,y)=" "xy\ +3y^{2}" "-2"



"f(1,2)=2" .......1)


"fx(1,2)=y=2" ......2)


"fy(1,2)\ =x+6y=13" .........3)


Putting the value of 1, 2 and 3 in equation B)...


"L(X,Y)=2+2(x-1)\ +13(y-2)=2x\ +13y\ -26"



"fxx(1,2)=0" ........4)


"fyy(1,2)=6" ..........5)


"fxy(1,2)=" "x+7y=15..........6)"



Putting 4, 5 and 6 and equation B in equation A , we get-



"Q(x,y)=2x\ +13y\ -26\ + 15(x-1)(x-2)\ +3(y-2)^{2}" which is required second taylor polynomial.